OpenStudy (anonymous):

oh man... i need help guys... derp Divide. x^2+2x+1 over x-2 over x^2-1 over x^2-4

OpenStudy (anonymous):

Just factor each expression and reduce common factors. \[=\frac{ (x+1)^{2} }{ (x-2)}\div \frac{ (x-1)(x+1) }{ (x-2)(x+2) }\] Now since we are dividing fractions, we have a rule that says you must multiply by the reciprocal of the second fraction. So, this leads to : \[\frac{ (x+1)^{2} }{ (x-2) }\times \frac{ (x-2)(x+2) }{ (x-1)x+1) }\] I think you should be able to finish it from there.